Solve: xp^2- (2x+3y)p + 6y=0 , p= dy/dx
1
Expert's answer
2020-11-05T10:32:33-0500

"xp^2-(2x+3y)p+6y=x(p^2-2p)-3y(p-2)=xp(p-2)-3y(p-2)"=

"=(xp-3y)(p-2)"=0

so we have two equations:


"x\frac{dy}{dx}-3y=0, \frac{dy}{dx}-2=0"


from the first equation:

"\frac{dy}{y}=3\frac{dx}{x}"

"\ln{y}=3\ln{x}+C"

"y=C'x^3" - the first solution


from the second equation:

"\frac{dy}{dx}=2"

"y=2x+C" - the second solution